The series combination of 5 F and 20 F capacitors is connected in parallel to 4 F capacitor. What is the total charge stored in the combination of capacitors when connected to 100V

1
Expert's answer
2022-03-20T18:48:05-0400

"C=\frac{5\cdot 20}{5+20}+4=8~F,"

"q=CU=8\cdot 100=800~C."